A traveller decides to go on vacation to Mexico. There are no direct flights from the​ traveller's city, and therefore her flight to Mexico will require a connection. The traveller is considering one of three different airlines to fly to​ Mexico: Airline​ A, Airline​ B, or Airline C. Because of the​ connection, the​ traveller's luggage is required to move from one airplane to​ another, as is the case with all connecting flights. After a​ Google-search, she has discovered that Airline A will misplace 12​% of all luggage it​ transfers, Airline B will misplace 4​% of all luggage it​ transfers, and Airline C will misplace 9​% of all luggage it handles. ​Historically, she has flown with Airline A 27​% of the​ time, with Airline B 48​% of the​ time, with the remaining percentage of her flights to Mexico being with Airline C. ​a) What is the probability that she arrives in​ Mexico, her luggage will arrive with her. That​ is, find the probability that this​ traveller's luggage will be not be misplaced. ​b) If her luggage was​ misplaced, what is the probability that she flew with Airline​ A?
Answer #1

a) P( she arrives in​ Mexico, her luggage will arrive with her) =P(travel by A)P(not missing luggage)+P(B)*P(not missing)+P(C)*P(not missing)=0.88*0.27+0.96*0.48+0.91*(1-0.27-0.48)=0.9259

b) P(lugguage misplaced) = 1-P(luggage arrive) = 1-0.9259=0.0741

P(flew with airline A|luggage misplaced) = P(airline A and luggage misplaced)/P(luggage misplaced) = 0.12*0.27/0.0741 = 0.4372

------------------------------

please revert if have any doubts